OpenStudy (anonymous):

find the missing probability? x 1 2 3 4 P(X) 0.30 0.40 0.15 ?

jimthompson5910 (jim_thompson5910):

Hint: in any valid probability distribution, all the probabilities must add to 1

OpenStudy (anonymous):

soo that means???

jimthompson5910 (jim_thompson5910):

0.3 + 0.4 + 0.15 + x = 1 solve for x

OpenStudy (anonymous):

idk what x is lol thats why i put it on here. i cant miss this one or i fail. sorry im not good with thid

jimthompson5910 (jim_thompson5910):

x is unknown and you don't know it until you solve for it

jimthompson5910 (jim_thompson5910):

0.3 + 0.4 + 0.15 + x = 1 0.85 + x = 1 x = 1 - 0.85 x = ???

jimthompson5910 (jim_thompson5910):

if we knew what x was, then there would be no need to ask the original question

OpenStudy (anonymous):

do i just minus 1 from 0.85?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

1 - 0.85 = ???

OpenStudy (anonymous):

i got 0.15

jimthompson5910 (jim_thompson5910):

same here

OpenStudy (anonymous):

so is that right?

jimthompson5910 (jim_thompson5910):

yes, that's the missing probability

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

basically you have 4 different possibilities possibility A has a 30% chance possibility B has a 40% chance possibility C has a 15% chance possibility D has a 15% chance (you just found this one) They add to 100% and there are no other possibilities

jimthompson5910 (jim_thompson5910):

and they must add to 100% if these are the only 4 possibilities